# The Belle II Collaboration

### v2

@@ Abstract
Despite clear observations of the gravitational effects of dark matter (DM) [1, 2], the mass of dark matter particles is unknown, and they have not yet been shown to interact with standard model (SM) particles. Direct detection experiments are sensitive to elastically scattering DM particles with masses typically in the GeV to TeV range [3, 4, 5, 6]. At colliders, searches for particles mediating interactions with dark sector particles are sensitive to mediator masses in the MeV to TeV range [7, 8]. These searches cover both promptly decaying and long-lived mediators decaying after a macroscopically large distance. A nonminimal class of models introduces inelastic DM (iDM), where DM couples inelastically to SM states, depending on the mass difference $\Delta m=m(\chi_{2})-m(\chi_{1})$ between the two DM mass eigenstates $\chi_{1}$ and $\chi_{2}$[9]. The simplest iDM models introduce the inelastic coupling via a massive dark photon $A^{\prime}$  that couples off diagonally to the two DM states. The $A^{\prime}$ kinetically mixes with SM photons via a mixing parameter $\epsilon$[10, 11, 12] and decays predominantly via $A^{\prime}\to\chi_{1}\chi_{2}$. A small $\Delta m$ or small coupling to the $A^{\prime}$  makes the heavier state $\chi_{2}$ long lived before it decays into a pair of SM particles and the lighter state $\chi_{1}$. The relic DM candidate $\chi_{1}$, which contributes to the DM abundance observed today, is stable and escapes detection [10]. These models can be extended to explain the mass splitting $\Delta m$ and the $A^{\prime}$  mass by introducing an additional dark Higgs boson $h^{\prime}$[13]. The $h^{\prime}$  would mix with the SM Higgs boson through a mixing angle $\theta$ [14, 15]. In total, the model has seven free parameters: the masses $m(h^{\prime}$), $m(A^{\prime}$), $m(\chi_{1}$), and the mass splitting $\Delta m$; the mixing angle $\theta$; the kinetic mixing parameter $\epsilon$; and the coupling $g_{D}=\sqrt{4\pi\alpha_{D}}$ between DM and the $A^{\prime}$. The coupling $k\approx g_{D}\Delta m/m(A^{\prime})$ between DM and the $h^{\prime}$  is fixed by the other parameters [13]. We restrict the search to parameter combinations that correspond to the perturbative regime and that evade existing constraints from observations of the cosmic microwave background by Planck [1] by requiring $m(h^{\prime})<m(\chi_{1})<m(A^{\prime})$. In this scenario, the DM relic density would be predominantly determined by the process $\chi_{1}\chi_{1}\to h^{\prime}h^{\prime}$. The $h^{\prime}$ lifetime increases for decreasing values of $\theta$, making the $h^{\prime}$ long-lived at small $\theta$.
@@
This model is already constrained by searches for $h^{\prime}$ or $A^{\prime}$ mediators without specific assumptions about an iDM model. Searches for scalars exclude $\sin\theta$ larger than $10^{-3}$ to $10^{-4}$ for $h^{\prime}$  masses up to about 5 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}  (see Ref. [16] for a review) for $m(h^{\prime})<2m(\chi_{1})$, while for higher $h^{\prime}$  masses the limits are considerably weaker. For the direct production of an $A^{\prime}$ through kinetic mixing with a photon and subsequent decay into iDM, the CMS experiment excludes $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ larger than $10^{-7}$ to $10^{-8}$ for $m(\chi_{1})\gtrsim 3\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$[17]. Reinterpretations [12, 18, 19] of searches for invisible decays of $A^{\prime}$  at BABAR [20], and of searches for long-lived $A^{\prime}$  decays at NuCal [21, 22], CHARM [23], and NA64 [24] exclude $y$ larger than about $10^{-9}$ for $m(\chi_{1})\lesssim 3\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ and $y$ larger than $10^{-12}$ below 1 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}.
@@
We use a $\mbox{365}~\mbox{\,fb}^{-1}$ data sample [25] collected at a center-of-mass (c.m.) energy of $\sqrt{s}=10.58\mathrm{\,Ge\kern-1.00006ptV}$ by the Belle II experiment [26] at the SuperKEKB $e^{+}e^{-}$ collider [27]. The beam energies are $7\mathrm{\,Ge\kern-1.00006ptV}$ for $e^{-}$ and $4\mathrm{\,Ge\kern-1.00006ptV}$ for $e^{+}$, resulting in a boost $\beta\gamma=0.28$ of the c.m. frame relative to the laboratory frame.
@@
We use simulated events to optimize the event selection, and to determine efficiencies and signal resolutions. Signal events are generated using a combination of MadGraph5@NLO [28] and EVTGEN [29] taking into account effects of initial state radiation (ISR) [30]. Furthermore, we consider effects from electromagnetic final state radiation in the decay of the $h^{\prime}$ using PHOTOS [31, 32]. To correct for efficiency differences caused by different beam background conditions, we generate signal simulations for a variety of different data-taking conditions using beam-induced backgrounds sampled from data overlaid with simulated signal events and find an approximately linear correlation between background level and signal efficiency. We use the efficiency obtained from a linear fit at the luminosity-weighted average beam background level of our dataset. Motivated by Ref. [13] and a previous search for a similar model [17], we consider values of $m(A^{\prime})=3\,m(\chi_{1})$ and $4\,m(\chi_{1})$, $\Delta m=0.2\,m(\chi_{1}),0.4\,m(\chi_{1})$ and $1.0\,m(\chi_{1})$, and $\alpha_{D}=0.1$ and $0.5$. For all possible combinations of these values, we generate events for $h^{\prime}$  masses between 0.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 3.0 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} in about 45 steps of varying size and various lifetimes $0.1<c\tau(h^{\prime})<10\,000$$\mathrm{\,cm}$} in steps that are approximately equidistant on a logarithmic scale; we generate events for $m(\chi_{1})$ between 0.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 3.0 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} in 30 steps of 0.1 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}, and various lifetimes $0.01<c\tau(\chi_{2})<1000$$\mathrm{\,cm}$} in variable steps. Since $m(A^{\prime})>m(\chi_{1})+m(\chi_{2})$, the $A^{\prime}$ in the $\chi_{2}$ decay is always off shell, while the $A^{\prime}$ produced in association with the $h^{\prime}$ can be either on shell or off shell with $A^{\prime}$ masses up to 12 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}. The lifetime of the $A^{\prime}$  is negligible for all values of $\epsilon$ probed in this analysis.
@@
We simulate the following background processes: $e^{+}e^{-}\to\Upsilon(4S)\to B\overline{B}$ with EvtGen; $e^{+}e^{-}\to q\bar{q}(\gamma)$, where $q=u,d,s,c$ with KKMC [33] interfaced with PYTHIA8 [34] and EvtGen; $e^{+}e^{-}\to\mu^{+}\mu^{-}(\gamma)$ and $e^{+}e^{-}\to\tau^{+}\tau^{-}(\gamma)$ with KKMC; $^{+}e^{-}\to e^{+}e^{-}e^{+}e^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\mu^{+}\mu^{-}$, $^{+}e^{-}\to\mu^{+}\mu^{-}\mu^{+}\mu^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\tau^{+}\tau^{-}$, and $e^{+}e^{-}\to\mu^{+}\mu^{-}\tau^{+}\tau^{-}$ with AAFH [35]; $e^{+}e^{-}\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ with KoralW [36]; $^{+}e^{-}\to e^{+}e^{-}\pi^{+}\pi^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}K^{+}K^{-}$, and $e^{+}e^{-}\to e^{+}e^{-}p\bar{p}$ with TREPS [37]; $e^{+}e^{-}\to e^{+}e^{-}(\gamma)$ and $e^{+}e^{-}\to\gamma\gamma(\gamma)$ with Babayaga.NLO [38]; $^{+}e^{-}\to K^{\rm\scriptstyle 0}_{\rm\scriptscriptstyle S}K^{\rm\scriptstyle 0}_{\rm\scriptscriptstyle L}\gamma, e+e$$^{+}e^{-}\to\pi^{+}\pi^{-}\gamma$, $e^{+}e^{-}\to K^{+}K^{-}\gamma$, and $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}\gamma$ with PHOKHARA [39]. Decays of $\tau$ leptons are simulated with TAUOLA [40] for KKMC, and using PYTHIA8 for all other event generators. The detector geometry and interactions of final-state particles with detector material are simulated using GEANT4 [41]. Both experimental and simulated events are reconstructed and analyzed using the Belle II software [42, 43]. To avoid experimenter’s bias, we examine the experimental data only after finalizing the analysis selection. All selection criteria are chosen by iteratively optimizing the figure of merit for a discovery with a significance of five standard deviations [44]. To avoid additional complexity of the analysis we chose a single set of selections for all model parameter combinations.
@@
PID information from all relevant subdetectors is combined to classify final states [45]. At least one track from the $h^{\prime}$  decay must have an extrapolated polar angle $37^{\circ}<\theta_{\rm{polar}}^{\rm{ext}}<122^{\circ}$, calculated by extrapolating the track helix to the KLM inner surface to ensure high muon identification efficiency. To further reduce the backgrounds in the final state with $h^{\prime}\to\pi^{+}\pi^{-}$ decays, all four tracks must be in the range $18^{\circ}<\theta_{\rm{polar}}^{\rm{ext}}<155^{\circ}$. To ensure high electron PID purity, we require the ratio between the energy deposition in the calorimeter and the momentum of the corresponding track to be larger than 0.6.
@@
We determine the signal mass resolution by fitting a double-sided Crystal Ball (DSCB) function [46, 47] to simulated $m(h^{\prime})$ distributions. The resolution $\sigma_{\mathrm{sig}}^{\mathrm{DSCB}}$ increases smoothly from about 1 $\mathrm{\,Me\kern-1.00006ptV\!/}c^{2}$} for a light $h^{\prime}$  to about 7 $\mathrm{\,Me\kern-1.00006ptV\!/}c^{2}$} for a heavy $h^{\prime}$  and depends only slightly on the $h^{\prime}$ lifetime or final state. Mass hypotheses that lack a simulation sample are interpolated from adjacent simulated samples.
@@
We evaluate systematic uncertainties affecting selection efficiency, integrated luminosity, the limited number of simulated events, and the background model. The dominant systematic uncertainties are associated with the signal efficiency, and depend on combinations of the $h^{\prime}$ and $\chi_{2}$ masses and lifetimes. Relative uncertainties are typically around 4% for most parameter configurations, but can reach 10% for the lightest $h^{\prime}$ masses and large displacements. For large displacements, the dominant systematic uncertainty on the signal efficiency arises from data-simulation differences in track finding for displaced tracks. We correct for this with an auxiliary measurement using $K_{S}^{0}$ decays from the process $D^{*+}\rightarrow D^{0}(\rightarrow K_{S}^{0}\pi^{+}\pi^{-})\pi^{+}$, and assign an uncertainty obtained by varying the nominal correction within the total uncertainty on the correction. For tracks close to the IP we instead correct for momentum-dependent efficiency differences between data and simulation, resulting in uncertainties at the level of 0.5% [48]. The uncertainties arising from PID are evaluated using the processes $^{+}e^{-}\to e^{+}e^{-}e^{+}e^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\mu^{+}\mu^{-}$, and $e^{+}e^{-}\to\mu^{+}\mu^{-}\gamma$, as well as decays of $K_{S}^{0}$ or $J/\Psi$. For electrons these uncertainties are typically at the level of 3%, while uncertainties for $\mu$, $\pi$, and $K$ are below 1%. We account for a lifetime-dependent effect on PID by introducing an additional systematic uncertainty, evaluated using $K_{S}^{0}$ and $\Lambda$ decays. For very displaced vertices, these uncertainties can reach up to 10%. The uncertainty on the luminosity is 0.47% [25]. The limited number of simulated events for each signal configuration introduces systematic uncertainties at the level of 1–2% for most parameter configurations but can reach up to 10% for very long lifetimes. We verify that our interpolation procedure between simulated mass points does not introduce a significant additional uncertainty. We estimate the uncertainty introduced by splitting the mass region in the $\pi^{+}\pi^{-}$ final state by varying the split point to 0.9 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 1.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}, respectively, and take the maximum deviation from the nominal background level as the uncertainty $\delta$.
@@
We find no events in the $\mu^{+}\mu^{-}$ final state, 8 events in the $\pi^{+}\pi^{-}$ final state, and one event in the $K^{+}K^{-}$ final state. The $M(\pi^{+}\pi^{-})$ distribution in the $h^{\prime}\to\pi^{+}\pi^{-}$ final state is shown in fig. 2, while the distributions for $h^{\prime}\to\mu^{+}\mu^{-}$ and $h^{\prime}\to K^{+}K^{-}$ are shown in the Supplemental Material [49]. The statistical model used to compute the signal significances and $p$ values is discussed in Appendix A. The largest local significance for the model-independent search is $2.9\,\sigma$, including systematic uncertainties, found near $m(h^{\prime})=0.531\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ for the $\pi^{+}\pi^{-}$ final state for a lifetime of $c\tau=1.0{\mathrm{\,cm}}$. Taking into account the look-elsewhere effect [50], this excess has a global significance of $1.1\,\sigma$.
@@
With the method described in Appendix B, we compute 95% Bayesian credibility level upper limits on $\sigma_{\mathrm{sig}}=\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)\times\mathcal{B}\left(h^{\prime}\to x^{+}x^{-}\right){}$ using the Bayesian Analysis Toolkit software package [51, 52]. The observed upper limits, including systematic uncertainties, are shown in the Supplemental Material [49]. Using a Jeffreys prior [53] would decrease the upper limits on $\sigma_{\mathrm{sig}}$ by up to 30% with respect to the uniform prior. The systematic uncertainties weaken the limits, with the largest increase of 2.5% occurring for heavy $h^{\prime}$ with small lifetimes.
@@
For the calculation of the model-dependent upper limits on $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$ we multiply the individual likelihoods weighted by the theoretical $h^{\prime}$ branching fractions from Ref. [54]. For each $h^{\prime}$ mass value, we determine the value of $\sin\theta$ such that the resulting predicted value of $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$, equals the 95% excluded $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$. To calculate the prediction, we fix $\sigma_{\mathrm{prod}}$ and the $\chi_{2}$ branching fractions to the theoretical values from Ref. [13] taking into account ISR. Figure 3 shows the observed upper limit on $\sin\theta$ for one specific choice of model parameters. Similarly, for each $m(\chi_{1}{})$, we determine the value of $y$ such that the resulting predicted value of $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$, equals the 95% excluded $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$. Figure 4 shows the observed upper limit on $y$ for a specific choice of model parameters. In general, $\sigma_{\mathrm{prod}}$ increases with $\epsilon^{2}$, the lifetime of the $h^{\prime}$ increases with $1/(\sin\theta)^{2}$, and the lifetime of the $\chi_{2}$ increases with $1/\epsilon^{2}$. Additional plots and detailed numerical results for many more parameter combinations can be found in the Supplemental Material [49].

### v3

@@ Abstract
Despite clear observations of the gravitational effects of dark matter (DM) [Planck:2018vyg, Arbey:2021gdg], the mass of dark matter particles is unknown, and they have not yet been shown to interact with standard model (SM) particles. Direct detection experiments are sensitive to elastically scattering DM particles with masses typically in the GeV to TeV range [LZ:2024zvo, XENON:2023cxc, PandaX:2024qfu, DarkSide-50:2022qzh]. At colliders, searches for particles mediating interactions with dark sector particles are sensitive to mediator masses in the MeV to TeV range [CMS:2024zqs, ATLAS:2024fdw]. These searches cover both promptly decaying and long-lived mediators decaying after a macroscopically large distance. A nonminimal class of models introduces inelastic DM (iDM), where DM couples inelastically to SM states, depending on the mass difference $\Delta m=m(\chi_{2})-m(\chi_{1})$ between the two DM mass eigenstates $\chi_{1}$ and $\chi_{2}$[Tucker-Smith:2001myb]. The simplest iDM models introduce the inelastic coupling via a massive dark photon $A^{\prime}$  that couples off diagonally to the two DM states. The $A^{\prime}$ kinetically mixes with SM photons via a mixing parameter $\epsilon$[Izaguirre:2015zva, Izaguirre:2017bqb, Duerr:2019dmv] and decays predominantly via $A^{\prime}\to\chi_{1}\chi_{2}$. A small $\Delta m$ or small coupling to the $A^{\prime}$  makes the heavier state $\chi_{2}$ long lived before it decays into a pair of SM particles and the lighter state $\chi_{1}$. The relic DM candidate $\chi_{1}$, which contributes to the DM abundance observed today, is stable and escapes detection [Izaguirre:2015zva]. These models can be extended to explain the mass splitting $\Delta m$ and the $A^{\prime}$  mass by introducing an additional dark Higgs boson $h^{\prime}$[Duerr:2020muu]. The $h^{\prime}$  would mix with the SM Higgs boson through a mixing angle $\theta$ [OConnell:2006rsp, Beacham:2019nyx]. In total, the model has seven free parameters: the masses $m(h^{\prime}$), $m(A^{\prime}$), $m(\chi_{1}$), and the mass splitting $\Delta m$; the mixing angle $\theta$; the kinetic mixing parameter $\epsilon$; and the coupling $g_{D}=\sqrt{4\pi\alpha_{D}}$ between DM and the $A^{\prime}$. The coupling $k\approx g_{D}\Delta m/m(A^{\prime})$ between DM and the $h^{\prime}$  is fixed by the other parameters [Duerr:2020muu]. We restrict the search to parameter combinations that correspond to the perturbative regime and that evade existing constraints from observations of the cosmic microwave background by Planck [Planck:2018vyg] by requiring $m(h^{\prime})<m(\chi_{1})<m(A^{\prime})$. In this scenario, the DM relic density would be predominantly determined by the process $\chi_{1}\chi_{1}\to h^{\prime}h^{\prime}$. The $h^{\prime}$ lifetime increases for decreasing values of $\theta$, making the $h^{\prime}$ long-lived at small $\theta$.
@@
This model is already constrained by searches for $h^{\prime}$ or $A^{\prime}$ mediators without specific assumptions about an iDM model. Searches for scalars exclude $\sin\theta$ larger than $10^{-3}$ to $10^{-4}$ for $h^{\prime}$  masses up to about 5 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}  (see Ref. [Ferber:2023iso] for a review) for $m(h^{\prime})<2m(\chi_{1})$, while for higher $h^{\prime}$  masses the limits are considerably weaker. For the direct production of an $A^{\prime}$ through kinetic mixing with a photon and subsequent decay into iDM, the CMS experiment excludes $y=\epsilon^{2}\alpha_{D}(m(\chi_{1})/m(A^{\prime}))^{4}$ larger than $10^{-7}$ to $10^{-8}$ for $m(\chi_{1})\gtrsim 3\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$[CMS:2023bay]. Reinterpretations [Duerr:2019dmv, Tsai:2019buq, Mongillo:2023hbs] of searches for invisible decays of $A^{\prime}$  at BABAR [BaBar:2017tiz], and of searches for long-lived $A^{\prime}$  decays at NuCal [Blumlein:2011mv, Blumlein:2013cua], CHARM [Gninenko:2012eq], and NA64 [NA64:2019auh] exclude $y$ larger than about $10^{-9}$ for $m(\chi_{1})\lesssim 3\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ and $y$ larger than $10^{-12}$ below 1 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}.
@@
We use a $\mbox{365}~\mbox{\,fb}^{-1}$ data sample [ref:lumi_paper] collected at a center-of-mass (c.m.) energy of $\sqrt{s}=10.58\mathrm{\,Ge\kern-1.00006ptV}$ by the Belle II experiment [Abe:2010gxa] at the SuperKEKB $e^{+}e^{-}$ collider [Akai:2018mbz]. The beam energies are $7\mathrm{\,Ge\kern-1.00006ptV}$ for $e^{-}$ and $4\mathrm{\,Ge\kern-1.00006ptV}$ for $e^{+}$, resulting in a boost $\beta\gamma=0.28$ of the c.m. frame relative to the laboratory frame.
@@
We use simulated events to optimize the event selection, and to determine efficiencies and signal resolutions. Signal events are generated using a combination of MadGraph5@NLO [Alwall:2014hca] and EVTGEN [Lange:2001uf] taking into account effects of initial state radiation (ISR) [Frixione:2021zdp]. Furthermore, we consider effects from electromagnetic final state radiation in the decay of the $h^{\prime}$ using PHOTOS [Barberio:1990ms, Barberio:1993qi]. To correct for efficiency differences caused by different beam background conditions, we generate signal simulations for a variety of different data-taking conditions using beam-induced backgrounds sampled from data overlaid with simulated signal events and find an approximately linear correlation between background level and signal efficiency. We use the efficiency obtained from a linear fit at the luminosity-weighted average beam background level of our dataset. Motivated by Ref. [Duerr:2020muu] and a previous search for a similar model [CMS:2023bay], we consider values of $m(A^{\prime})=3\,m(\chi_{1})$ and $4\,m(\chi_{1})$, $\Delta m=0.2\,m(\chi_{1}),0.4\,m(\chi_{1})$ and $1.0\,m(\chi_{1})$, and $\alpha_{D}=0.1$ and $0.5$. For all possible combinations of these values, we generate events for $h^{\prime}$  masses between 0.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 3.0 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} in about 45 steps of varying size and various lifetimes $0.1<c\tau(h^{\prime})<10\,000$$\mathrm{\,cm}$} in steps that are approximately equidistant on a logarithmic scale; we generate events for $m(\chi_{1})$ between 0.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 3.0 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} in 30 steps of 0.1 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}, and various lifetimes $0.01<c\tau(\chi_{2})<1000$$\mathrm{\,cm}$} in variable steps. Since $m(A^{\prime})>m(\chi_{1})+m(\chi_{2})$, the $A^{\prime}$ in the $\chi_{2}$ decay is always off shell, while the $A^{\prime}$ produced in association with the $h^{\prime}$ can be either on shell or off shell with $A^{\prime}$ masses up to 12 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}. The lifetime of the $A^{\prime}$  is negligible for all values of $\epsilon$ probed in this analysis.
@@
We simulate the following background processes: $e^{+}e^{-}\to\Upsilon(4S)\to B\overline{B}$ with EvtGen; $e^{+}e^{-}\to q\bar{q}(\gamma)$, where $q=u,d,s,c$ with KKMC [Jadach:1999vf] interfaced with PYTHIA8 [Sjostrand:2014zea] and EvtGen; $e^{+}e^{-}\to\mu^{+}\mu^{-}(\gamma)$ and $e^{+}e^{-}\to\tau^{+}\tau^{-}(\gamma)$ with KKMC; $^{+}e^{-}\to e^{+}e^{-}e^{+}e^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\mu^{+}\mu^{-}$, $^{+}e^{-}\to\mu^{+}\mu^{-}\mu^{+}\mu^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\tau^{+}\tau^{-}$, and $e^{+}e^{-}\to\mu^{+}\mu^{-}\tau^{+}\tau^{-}$ with AAFH [Berends:1984gf]; $e^{+}e^{-}\to\tau^{+}\tau^{-}\tau^{+}\tau^{-}$ with KoralW [Jadach:1998gi]; $^{+}e^{-}\to e^{+}e^{-}\pi^{+}\pi^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}K^{+}K^{-}$, and $e^{+}e^{-}\to e^{+}e^{-}p\bar{p}$ with TREPS [Uehara:1996bgt]; $e^{+}e^{-}\to e^{+}e^{-}(\gamma)$ and $e^{+}e^{-}\to\gamma\gamma(\gamma)$ with Babayaga.NLO [Balossini:2008xr]; $^{+}e^{-}\to K^{\rm\scriptstyle 0}_{\rm\scriptscriptstyle S}K^{\rm\scriptstyle 0}_{\rm\scriptscriptstyle L}\gamma, e+e$$^{+}e^{-}\to\pi^{+}\pi^{-}\gamma$, $e^{+}e^{-}\to K^{+}K^{-}\gamma$, and $e^{+}e^{-}\to\pi^{+}\pi^{-}\pi^{0}\gamma$ with PHOKHARA [Campanario:2019mjh]. Decays of $\tau$ leptons are simulated with TAUOLA [Jadach:1990mz] for KKMC, and using PYTHIA8 for all other event generators. The detector geometry and interactions of final-state particles with detector material are simulated using GEANT4 [Agostinelli:2002hh]. Both experimental and simulated events are reconstructed and analyzed using the Belle II software [Kuhr:2018lps, basf2-zenodo]. To avoid experimenter’s bias, we examine the experimental data only after finalizing the analysis selection. All selection criteria are chosen by iteratively optimizing the figure of merit for a discovery with a significance of five standard deviations [Punzi:2003bu]. To avoid additional complexity of the analysis we chose a single set of selections for all model parameter combinations.
@@
PID information from all relevant subdetectors is combined to classify final states [Kou:2018nap]. At least one track from the $h^{\prime}$  decay must have an extrapolated polar angle $37^{\circ}<\theta_{\rm{polar}}^{\rm{ext}}<122^{\circ}$, calculated by extrapolating the track helix to the KLM inner surface to ensure high muon identification efficiency. To further reduce the backgrounds in the final state with $h^{\prime}\to\pi^{+}\pi^{-}$ decays, all four tracks must be in the range $18^{\circ}<\theta_{\rm{polar}}^{\rm{ext}}<155^{\circ}$. To ensure high electron PID purity, we require the ratio between the energy deposition in the calorimeter and the momentum of the corresponding track to be larger than 0.6.
@@
We determine the signal mass resolution by fitting a double-sided Crystal Ball (DSCB) function [Gaiser:Phd, Skwarnicki:1986xj] to simulated $m(h^{\prime})$ distributions. The resolution $\sigma_{\mathrm{sig}}^{\mathrm{DSCB}}$ increases smoothly from about 1 $\mathrm{\,Me\kern-1.00006ptV\!/}c^{2}$} for a light $h^{\prime}$  to about 7 $\mathrm{\,Me\kern-1.00006ptV\!/}c^{2}$} for a heavy $h^{\prime}$  and depends only slightly on the $h^{\prime}$ lifetime or final state. Mass hypotheses that lack a simulation sample are interpolated from adjacent simulated samples.
@@
We evaluate systematic uncertainties affecting selection efficiency, integrated luminosity, the limited number of simulated events, and the background model. The dominant systematic uncertainties are associated with the signal efficiency, and depend on combinations of the $h^{\prime}$ and $\chi_{2}$ masses and lifetimes. Relative uncertainties are typically around 4% for most parameter configurations, but can reach 10% for the lightest $h^{\prime}$ masses and large displacements. For large displacements, the dominant systematic uncertainty on the signal efficiency arises from data-simulation differences in track finding for displaced tracks. We correct for this with an auxiliary measurement using $K_{S}^{0}$ decays from the process $D^{*+}\rightarrow D^{0}(\rightarrow K_{S}^{0}\pi^{+}\pi^{-})\pi^{+}$, and assign an uncertainty obtained by varying the nominal correction within the total uncertainty on the correction. For tracks close to the IP we instead correct for momentum-dependent efficiency differences between data and simulation, resulting in uncertainties at the level of 0.5% [PhysRevD.111.032012]. The uncertainties arising from PID are evaluated using the processes $^{+}e^{-}\to e^{+}e^{-}e^{+}e^{-}, e+e$$^{+}e^{-}\to e^{+}e^{-}\mu^{+}\mu^{-}$, and $e^{+}e^{-}\to\mu^{+}\mu^{-}\gamma$, as well as decays of $K_{S}^{0}$ or $J/\Psi$. For electrons these uncertainties are typically at the level of 3%, while uncertainties for $\mu$, $\pi$, and $K$ are below 1%. We account for a lifetime-dependent effect on PID by introducing an additional systematic uncertainty, evaluated using $K_{S}^{0}$ and $\Lambda$ decays. For very displaced vertices, these uncertainties can reach up to 10%. The uncertainty on the luminosity is 0.47% [ref:lumi_paper]. The limited number of simulated events for each signal configuration introduces systematic uncertainties at the level of 1–2% for most parameter configurations but can reach up to 10% for very long lifetimes. We verify that our interpolation procedure between simulated mass points does not introduce a significant additional uncertainty. We estimate the uncertainty introduced by splitting the mass region in the $\pi^{+}\pi^{-}$ final state by varying the split point to 0.9 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$} and 1.2 $\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}$}, respectively, and take the maximum deviation from the nominal background level as the uncertainty $\delta$.
@@
We find no events in the $\mu^{+}\mu^{-}$ final state, 8 events in the $\pi^{+}\pi^{-}$ final state, and one event in the $K^{+}K^{-}$ final state. The $M(\pi^{+}\pi^{-})$ distribution in the $h^{\prime}\to\pi^{+}\pi^{-}$ final state is shown in fig. 2, while the distributions for $h^{\prime}\to\mu^{+}\mu^{-}$ and $h^{\prime}\to K^{+}K^{-}$ are shown in the Supplemental Material [aux:2024]. The statistical model used to compute the signal significances and $p$ values is discussed in Appendix A. The largest local significance for the model-independent search is $2.9\,\sigma$, including systematic uncertainties, found near $m(h^{\prime})=0.531\,{\mathrm{\,Ge\kern-1.00006ptV\!/}c^{2}}$ for the $\pi^{+}\pi^{-}$ final state for a lifetime of $c\tau=1.0{\mathrm{\,cm}}$. Taking into account the look-elsewhere effect [Gross:2010qma], this excess has a global significance of $1.1\,\sigma$.
@@
With the method described in Appendix B, we compute 95% Bayesian credibility level upper limits on $\sigma_{\mathrm{sig}}=\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)\times\mathcal{B}\left(h^{\prime}\to x^{+}x^{-}\right){}$ using the Bayesian Analysis Toolkit software package [Caldwell:2008fw, bat-zenodo]. The observed upper limits, including systematic uncertainties, are shown in the Supplemental Material [aux:2024]. Using a Jeffreys prior [Jeffreys:1946] would decrease the upper limits on $\sigma_{\mathrm{sig}}$ by up to 30% with respect to the uniform prior. The systematic uncertainties weaken the limits, with the largest increase of 2.5% occurring for heavy $h^{\prime}$ with small lifetimes.
@@
For the calculation of the model-dependent upper limits on $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$ we multiply the individual likelihoods weighted by the theoretical $h^{\prime}$ branching fractions from Ref. [Winkler:2018qyg]. For each $h^{\prime}$ mass value, we determine the value of $\sin\theta$ such that the resulting predicted value of $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$, equals the 95% excluded $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$. To calculate the prediction, we fix $\sigma_{\mathrm{prod}}$ and the $\chi_{2}$ branching fractions to the theoretical values from Ref. [Duerr:2020muu] taking into account ISR. Figure 3 shows the observed upper limit on $\sin\theta$ for one specific choice of model parameters. Similarly, for each $m(\chi_{1}{})$, we determine the value of $y$ such that the resulting predicted value of $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$, equals the 95% excluded $\sigma_{\mathrm{prod}}\times\mathcal{B}\left(\chi_{2}\to\chi_{1}e^{+}e^{-}\right)$. Figure 4 shows the observed upper limit on $y$ for a specific choice of model parameters. In general, $\sigma_{\mathrm{prod}}$ increases with $\epsilon^{2}$, the lifetime of the $h^{\prime}$ increases with $1/(\sin\theta)^{2}$, and the lifetime of the $\chi_{2}$ increases with $1/\epsilon^{2}$. Additional plots and detailed numerical results for many more parameter combinations can be found in the Supplemental Material [aux:2024].